Using the income statement for Times Mirror and Glass Co., compute the following ratios TIMES MIRROR AND GLASS Co Income Statement Sales Cost of goods sold Gross profit Selling and administrative expense Lease expense Operating profit* Interest expense Earnings before taxes Taxes (30%) Earnings after taxes *Equals income before interest and taxes $ 220,000 131,000 $ 89,000 44,500 14,200 $ 30,300 10,400 $ 19,900 7,960 $ 11,940 a.Compute the interest coverage ratio. (Round your answer to 2 decimal places.) Interest coverage times b.Compute the fixed charge coverage ratio. (Round your answer to 2 decimal places.) Fixed charge coverage timesThe total assets for this company equal $215,000. Set up the equation for the Du Pont system of ratio analysis. с.Compute the profit margin ratio. (Input your answer as a percent rounded to 2 decimal places.) Profit margin d.Compute the total asset turnover ratio. (Round your answer to 2 decimal places.) Total asset turnover imes e. Compute the return on assets (investment). (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.) Return on assets
Answer #1

Please provide rating...............

Answer a) Interest coverage ratio = EBIT/Interest expenses
Interest coverage ratio = =30300/10400
         2.91
Answer b)
Fixed charge coverage = Income before fixed charges and interest/(Fixed charges+interest expense)
Fixed charge coverage = =(30300+14200)/(10400+14200)
         1.81
Answer c)
Profit margin ratio=(Net income/Sales)*100
Profit margin ratio = =(11940/220000)
5.43%
Answer d)
Total asset turnover ratio = Sales/Total asset
220000/215000
         1.02
Answer e) Return on asset = Net income/Total asset
=11940/215000
5.55%